# Investigation of Holographic Beamforming via Dynamic Metasurface Antennas in QoS Guaranteed Power Efficient Networks

### v2

@@ II SYSTEM MODEL & PROBLEM FORMULATION
We consider a multi-user downlink MISO [^1] We use the term MISO since receiver users have a single antenna. However, DMA-aided MISO networks are often termed MIMO, as DMA act as HMIMO surfaces with MIMO-like two-dimensional precoding although a single antenna receiver is considered [4]. system, where the base station (BS) is equipped with either FD or DMA-aided beamforming architectures. In this system, the BS serves $K$ users, each equipped with a single antenna and requesting a certain level of SINR, in a generic Line-of-Sight (LoS) channel model that covers both far-field and near-field users. Moreover, perfect knowledge of the channel is assumed.
@@ II-C General Problem Formulation
Given that optimization aim of the downlink beamforming problem is to minimize the total transmit power under the constraints of given sets of SINR for different users, the problem can be expressed as
@@ III-B Optimizing the DMA Weights
Finally, solving the individual problems (12) and (18) in an alternating way effectively leads to the convergence of the digital precoder vectors toward the optimal solution, while optimizing the DMA weights toward suboptimal solutions. This approach is encapsulated in the proposed algorithm, which is detailed in Algorithm 1. The worst-case complexity of the proposed algorithm can be determined in reference to the complexity of SDP using an interior point method, which is given with $O(\max\{m,n\}^{4}n^{\frac{1}{2}}\log(1/\epsilon))$, where $m$ is the number of constraints, $n$ denotes the number of optimization variables, and $\epsilon$ represents the accuracy [9]. Taking into account the Hermitian structure of the optimization variables $\mathbf{W}$ and $\mathbf{\tilde{Q}}$ in problems (12) and (18), the number of optimization variables can be determined with $n=\frac{N_{\text{r}}(N_{\text{r}}+1)}{2}$ for (12) and $n=\frac{N(N+1)}{2}$ for (18). Therefore, the overall complexity remains polynomial in both the problem size $n$ and the number of constraints $m$, ensuring computational feasibility .
III-C Discussion
When comparing the optimization problems for FD-based architecture (7) and DMA-based architecture (9), DMA has limitations in terms of reduced DoF due to the Lorentzian constraint in (9c) and the reduced number of optimization variables. The fewer optimization variables result from the reduced number of RF chains in DMA compared to FD, which helps lower transmitter power consumption. However, the reduced DoF may degrade beamforming performance and increase the required transmission power. This study focuses on quantifying the performance gap caused by these DoF limitations in DMA. We assess their impact on beamforming through numerical comparisons of optimization problems. The power consumption of the RF chain will be examined in future work. For comparison purposes, we consider two optimization problems OP1 and OP2. When the problem given in (8) is solved with the same antenna spacing of DMA configuration, the optimization problem is called OP1.
@@ III-C Discussion
Note that, OP1 corresponds to the solution of the FD-based architecture when $d_{x}=\lambda/2$. The number of optimization variables in OP1 is given by $\text{nDoF}=N_{\text{r}}N_{\text{c}}M$, whereas for DMA, it is $N_{\text{r}}(M+N_{\text{c}})$ (from (12) and (18)). This suggests that as the number of users increase, DMA will leverage less DoF with respect to OP1, leading to a performance gap in terms of beamforming performance. Furthermore, we introduce OP2, where the DMA optimization problem is solved with unrestricted weights to further analyze the impact of Lorentzian constraints on DoF reduction in DMA. Numerical comparisons with OP1 and OP2 will illustrate the impact of DoF limitations and Lorentzian constraints on beamforming performance.
@@ IV NUMERICAL RESULTS
In the numerical results, we consider the simulation setup that a uniform planar array is placed in the xy-plane, and the users are randomly distributed over 200222The near-field channel model used in this paper and in the related literature is deterministic, excluding randomness like Rayleigh fading. Prior studies, e.g., [4, 5], evaluate DMA performance for limited user positions with up to 30 realizations. Here, Monte Carlo simulations are increased to 200 to assess the transmitter performance across a region with randomly located users. realizations within the half-circles with radius $r$ lie in the xz-plane, covering both near-zone ($0.1d_{\text{F}}<r<1d_{\text{F}}$) and far-zone regions ($1d_{\text{F}}<r<5d_{\text{F}}$). Here, $d_{\text{F}}$ is the Fraunhofer distance, i.e., $d_{\text{F}}\triangleq\frac{2D^{2}}{\lambda}$ for antenna aperture length of $D$ and wavelength ${\lambda}$. Throughout the experimental study, the frequency is set to $f=28\,\text{GHz}$, and the aperture size is consistently fixed at $10\,\text{cm}\times 10\,\text{cm}$ for all structures. Optimizations are performed to guarantee the minimum achievable rate for each user ($R_{\text{min},k}$), which can be related to the SINR conditions in optimization problems by $\delta_{k}\triangleq(2^{R_{\text{min},k}}-1)$, where $R_{\text{min},k}$ is the same for all $k$. The noise power is set to $\sigma_{k}^{2}=-114\,\text{dBm}$. For DMA, flat frequency characteristics are achieved by assuming the $\mathbf{H}$ matrix is an identity matrix. For all structures, the separation between rows of the UPA is kept constant at $d_{y}=\lambda/2$, where $N_{\text{r}}=\left\lfloor 2D/\lambda\right\rfloor$. The separation between columns of the UPA, denoted by $d_{x}$, is varied for the case of DMA, while it is fixed at $d_{x}=\lambda/2$ for the FD-based architecture as a benchmark.

### v3

@@ II SYSTEM MODEL & PROBLEM FORMULATION
Figure 1: System configurations for different architectures: (a) FD, (b) DMA.
We consider a multi-user downlink MISO [^1] We use the term MISO since receiver users have a single antenna. However, DMA-aided MISO networks are often termed MIMO, as DMA act as HMIMO surfaces with MIMO-like two-dimensional precoding although a single antenna receiver is considered [4]. system, where the base station (BS) is equipped with either FD-aided (Fig. 1(a)) or DMA-aided (Fig. 1(b)) beamforming architectures. In this system, the BS serves $K$ users, each equipped with a single antenna and requesting a certain level of SINR, in a generic Line-of-Sight (LoS) channel model that covers both far-field and near-field users. Moreover, perfect knowledge of the channel is assumed.
@@ II-C General Problem Formulation
Given that optimization aim of the downlink beamforming problem is to minimize the total transmit power under the constraints of given sets of SINR for different users, the problem for generic FD case can be expressed as
@@ III-B Optimizing the DMA Weights
Finally, solving the individual problems (12) and (18) in an alternating way effectively leads to the convergence of the digital precoder vectors toward the optimal solution, while optimizing the DMA weights toward suboptimal solutions. This approach is encapsulated in the proposed algorithm, which is detailed in Algorithm 1. The worst-case complexity of the proposed algorithm can be determined in reference to the complexity of SDP using an interior point method, which is given with $O(\max\{m,n\}^{4}n^{\frac{1}{2}}\log(1/\epsilon))$, where $m$ is the number of constraints, $n$ denotes the number of optimization variables, and $\epsilon$ represents the accuracy [9]. In this, the number of optimization variables can be determined with $n=N_{\text{r}}$ for SDP problem in (12) and $n=N$ for SDP problem in (18), where the number of constraints $m$ in both SDP problems is equal to the number of users ($K$). Therefore, the overall complexity remains polynomial in both the problem size $n$, and the number of constraints $m$, ensuring computational feasibility.
@@
III-C Discussion
When comparing the optimization problems for FD-based architecture (7) and DMA-based architecture (9), DMA has limitations in terms of reduced DoF due to the Lorentzian constraint in (9c) and the reduced number of optimization variables. The fewer optimization variables result from the reduced number of RF chains in DMA compared to FD, which helps lower transmitter power consumption. However, the reduced DoF may degrade beamforming performance and increase the required transmission power. This study focuses on quantifying the performance gap caused by these DoF limitations in DMA. We assess their impact on beamforming through numerical comparisons of optimization problems. The power consumption of the RF chain will be examined in future work. For comparison purposes, we consider two optimization problems OP1 and OP2. When the problem given in (8) is solved with the same antenna spacing of DMA configuration, the optimization problem is called OP1. Note that, OP1 corresponds to the solution of the FD-based architecture when $d_{x}=\lambda/2$. The number of optimization variables in OP1 is given by $\text{nDoF}=N_{\text{r}}N_{\text{c}}M$, whereas for DMA, it is $N_{\text{r}}(M+N_{\text{c}})$ (from (12) and (18)). This suggests that as the number of users increase, DMA will leverage less DoF with respect to OP1, leading to a performance gap in terms of beamforming performance. Furthermore, we introduce OP2, where the DMA optimization problem is solved with unrestricted weights to further analyze the impact of Lorentzian constraints on DoF reduction in DMA. Numerical comparisons with OP1 and OP2 will illustrate the impact of DoF limitations and Lorentzian constraints on beamforming performance.
@@ IV NUMERICAL RESULTS
In the numerical results, we consider the simulation setup that a uniform planar array is placed in the xy-plane, and the users are randomly distributed over 200222The near-field channel model used in this paper and in the related literature is deterministic, excluding randomness like Rayleigh fading. Prior studies, e.g., [4, 5], evaluate DMA performance for limited user positions with up to 30 realizations. Here, Monte Carlo simulations are increased to 200 to assess the transmitter performance across a region with randomly located users. realizations within the half-circles with radius $r$ lie in the xz-plane, covering both near-zone ($0.1d_{\text{F}}<r<1d_{\text{F}}$) and far-zone regions ($1d_{\text{F}}<r<5d_{\text{F}}$). Here, $d_{\text{F}}$ is the Fraunhofer distance, i.e., $d_{\text{F}}\triangleq\frac{2D^{2}}{\lambda}$ for antenna aperture length of $D$ and wavelength ${\lambda}$. Throughout the experimental study, the frequency is set to $f=28\,\text{GHz}$, and the aperture size is consistently fixed at $10\,\text{cm}\times 10\,\text{cm}$ for all structures. Optimizations are performed to guarantee the minimum achievable rate for each user ($R_{\text{min},k}$), which can be related to the SINR conditions in optimization problems by $\delta_{k}\triangleq(2^{R_{\text{min},k}}-1)$, where $R_{\text{min},k}$ is the same for all $k$. The noise power is set to $\sigma_{k}^{2}=-114\,\text{dBm}$.
For DMA, flat frequency characteristics are achieved by assuming the $\mathbf{H}$ matrix is an identity matrix. For all structures, the separation between rows of the UPA is kept constant at $d_{y}=\lambda/2$, where $N_{\text{r}}=\left\lfloor 2D/\lambda\right\rfloor$. The separation between columns of the UPA, denoted by $d_{x}$, is varied for the case of DMA, while it is fixed at $d_{x}=\lambda/2$ for the FD-based architecture as a benchmark.